Testbank, Question 18

Montreal Growers Inc. issued 1 million preferred shares at a par value of $20 and the dividend rate is 10%. If the risk-free rate is 4% and the risk premium is 3%, what is the preferred share price?

$50.00

$28.57

$35.00

$14.00
Answer #1

Annual Dividend = 20*10% = $ 2.00 0.04+0.03 = 0.07 Price Annual Dividend/Required Return 2/0.07 $ 28.57